An intensity of a sound wave is measured to be 1.87×104 W/m2 at position A, 120 m away from the source of this sound. A worker at position B, 308 m away from the source hears the sound.







What is the sound level 120 m way from the source?

1
Expert's answer
2021-11-17T10:37:18-0500

Given:

"I_1=1.87*10^4\:\rm W/m^2"

"r_1=120\:\rm m"

"r_2=308\:\rm m"


The intensity is inversialy proportional to the square of distance. Hence


"I_2=I_1\frac{r_1^2}{r_2^2}=1.87*10^4*\frac{120^2}{308^2}=2839\:\rm W"